# A Matrix Model Proposal for Quantum Gravity and the Quantum Mechanics of Black Holes

### v2

@@ 5 Discussion
In this paper we have proposed a theory of quantum space and gravity as a large $N$ quantum mechanics of non-abelian bosonic and fermionic coordinates. The quantum mechanics has static solution whose bosonic part is given by a fuzzy sphere. Over the fuzzy sphere geometry, the fermionic part of the theory is given by a collection of fermionic oscillators all with the same frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This half-filled fuzzy sphere observes the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere radius is identified with the horizon size. Moreover, the coarse graining of the set of quantum states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. As a result, we propose that a quantum black hole is described by a half-filled fuzzy sphere in our theory.
We have also considered interaction between these fuzzy spheres by including the quantum mechanical perturbation arisen from a separation of them in the matrix space. We find that the interaction energy between the fuzzy spheres has a dependence on the masses and Newton constant exactly as in the Newton gravity. We conjecture that when the full large $N$ resummation of all the higher order perturbation effects including the off-diagonal block terms is performed, the exact result will capture the correct gravitational interaction between the static quantum black holes, with the Newton gravity reproduced in the large distance limit. The obtained modified gravity may be relevant for the dark matter and dark energy problem.
In this paper we have considered static solutions. It would be interesting to consider non-static solution in order to capture the relativistic effects. Our proposed quantum mechanics is for quantum space. Rotational black hole should be included in our model. It is interesting to understand whether and how matter would arise in our quantum mechanics. In particular, how gauge interaction can be incorporated and how a charged back hole can be obtained.
We note that the change of coordinates $x^{\mu}\to x^{\prime}{}^{\mu}(x)$ in general relativity is replaced in our description by the unitary transformation
@@ Acknowledgments
The support of this work by NCTS and the grant 110-2112-M-007-015-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.
@@ Appendix A Ensemble average of 1st order perturbation
In this appendix, we use perturbation theory to compute the average correction to the energy of the system due to the interaction term between the center coordinates of the fuzzy sphere with the fermions.
@@
$h_{0}=\psi^{\dagger}\sigma^{a}X^{0a}\psi,$ (A.1)
where $X^{0a}$ is given by a fuzzy sphere. In the large $N$ limit, $h_{0}$ can be written as a sum of fermionic oscillators with the same frequency $\omega$,
$h_{0}=\frac{N}{2}\left(\sum_{p,k=1}^{N}(\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p})-N^{2}\right),$ (A.2)
where the fermionic oscillators
$\xi_{k}^{p}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi_{k}^{p\dagger}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta}$ (A.3)
are defined using the eigenvectors ${\cal U}^{p},{\cal V}^{p}$ of the matrix $K:=\sigma^{a}X^{0a}$:
@@
They satisfy the orthonormality condition
@@
$h_{0}=n,$ (A.8)
where $n:=r+s-N^{2}=-N^{2},\cdots,0,\cdots N^{2}$ is the energy level within the Fermi sea.
We are interested in the correction to the energy state of $h_{0}$ due to term of the form
$h^{\prime}:=\sum_{m,n=1}^{N}\psi^{\dagger}_{mn}\sigma^{a}x^{a}\psi_{mn}.$ (A.9)
We will treat $h^{\prime}$ as a perturbation. The unperturbed level $n$ has a degeneracy of $\Omega_{n}=\binom{2N^{2}}{n+N^{2}}$. Denoting the set of unperturbed eigenstates by
@@
where $|i\rangle$ are those states (A.7) with $r+s=n+N^{2}$. With the perturbation $h^{\prime}$ turns on, some of these states will be lifted. Rather then trying to determine the splitting $\lambda_{p}$ for each individual states, it is more meaningful to compute the average correction to the level $n$ eigenvalue,
$\overline{h^{\prime}}:=\frac{1}{\Omega_{n}}\sum_{p=1}^{\Omega_{n}}\lambda_{p}=\frac{1}{\Omega_{n}}\sum_{|i\rangle\in\mathbb{V}_{n}}\langle i|h^{\prime}|i\rangle.$ (A.11)
We claim that $\overline{h^{\prime}}=0$. To see this, let us use
@@
$h^{\prime}=P^{pq}\xi^{p\dagger}_{k}\xi^{q}_{k}+Q^{pq}\chi^{p}_{k}\chi^{q\dagger}_{k}+T^{pq}\xi^{p\dagger}_{k}\chi^{q\dagger}_{k}+{\rm h.c.}$ (A.13)
@@
$P^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal U}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a},\quad Q^{pq}:={\cal V}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a},\quad$ $T^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a}.$ (A.14)
It is easy to see that the $T$ terms do not contribute to the matrix element and so
$\langle i|h^{\prime}|i\rangle=N\binom{N^{2}-1}{r}\binom{N^{2}}{s}\times\mbox{tr}P-N\binom{N^{2}-1}{s}\binom{N^{2}}{r}\times\mbox{tr}Q$ (A.15)
for any state $|i\rangle$ of the form (A.7) with $r$ and $s$ fixed. Using $\sum_{r}\binom{N^{2}-1}{r}\binom{N^{2}}{N^{2}-r}=\Omega_{0}/2$,the sum over all states of form (A.7) with a fixed level $n=0$ is given by
@@
$\overline{h^{\prime}}=\mbox{tr}(K\sigma^{a}X^{0a}).$ (A.17)
This vanishes since $\mbox{tr}(X^{0a})=0$ and hence our claim that $\overline{h^{\prime}}=0$.

### v3

@@ 6 Discussion
In this paper we have proposed a model of quantum space and gravity as a large $N$ quantum mechanics of non-abelian bosonic and fermionic coordinates. The quantum mechanics has static solution whose bosonic part is given by a fuzzy sphere. Over the fuzzy sphere geometry, the fermionic part of the theory is given by a collection of fermionic oscillators all with the same frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This half-filled fuzzy sphere observes the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere radius is identified with the horizon size. Moreover, the coarse graining of the set of quantum states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. As a result, we propose that a quantum black hole is described by a half-filled fuzzy sphere in our model.
We have also considered interaction between these fuzzy spheres by including the quantum mechanical perturbation arisen from a separation of them in the matrix space. We find that the interaction energy between the fuzzy spheres has a dependence on the masses and Newton constant exactly as in the Newton gravity. We conjecture that when the full large $N$ resummation of all the higher order quantum effects including the off-diagonal block terms is performed, the exact result will capture the correct gravitational interaction between the static quantum black holes, with the Newton gravity reproduced in the large distance limit. The obtained modified gravity may be relevant for the dark matter and dark energy problem.
In this paper we have considered static solutions. It would be interesting to consider non-static solution in order to capture the relativistic effects. Our proposed quantum mechanics is for quantum space. Rotational black hole should be included in our model. It is interesting to understand whether and how matter would arise in our quantum mechanics. In particular, how gauge interaction can be incorporated and how a charged back hole can be obtained. It is also interesting to consider solution with nontrivial fermionic background so as to arrive at a flat solution.
[…]
@@ Acknowledgments
We thank Andrew Cohen, Pei-Ming Ho, Hikaru Kawai and Harold Steinacker for interesting discussion. The support of this work by NCTS and the grants 110-2112-M-007-015-MY3 and 113-2112-M-007-039-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.
@@ Appendix A Ensemble average of 1st order perturbation
In this appendix, we use perturbation theory to compute the average correction to the energy of the fuzzy spheres system due to the fermionic term of the Hamiltonian.
@@
$h_{0}=\mathop{:}\psi^{\dagger}\sigma^{a}X^{0a}\psi\mathop{:},$ (A.1)
where $X^{0a}$ is given by a fuzzy sphere. In the large $N$ limit, the matrix $\sigma^{a}X^{0a}$ admits the eigenvectors ${\cal U}^{p},{\cal V}^{p}$ that corresponds to positive and negative eigenvalues N/2,$N/2$,-N/2 respectively. One has the eigen-decomposition
@@
The eigenvectors ${\cal U}^{p},{\cal V}^{p}$ satisfy the orthonormality condition
@@
Introducing the fermionic oscillators
$\xi_{k}^{p}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi_{k}^{p\dagger}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta},$ (A.5)
where the normal ordering is defined with respect to, then
$h_{0}=\frac{N}{2}\sum_{p,k=1}^{N}(\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p}).$ (A.6)
@@
$h_{0}=n+N^{2},$ (A.8)
where $n:=r+s-N^{2}=-N^{2},\cdots,0,\cdots N^{2}$ is the energy level within the Fermi sea. The energy level $n$ has a degeneracy of $\Omega_{n}=\binom{2N^{2}}{n+N^{2}}$. Denoting this set of states by
@@
where $|i\rangle$ are those states (A.7) with $r+s=n+N^{2}$. In the main body of the text, we have identified the set $\Omega_{0}$ of states with the microstates of the Schwarzschild black hole. Now let us consider the Hamiltonian $h=h_{0}+h^{\prime}$ with an operator $h^{\prime}$ of the form
$h^{\prime}:=\sum_{m,n=1}^{N}\mathop{:}\psi^{\dagger}_{mn}\tilde{h}^{\prime}\psi_{mn}\mathop{:}.$ (A.10)
We are interested in the change of energy of the fuzzy spheres system.
For generality, we will treat $h^{\prime}$ as a perturbation. Obviously, each of the microstates $|i\rangle\in\mathbb{V}_{0}$ will receive a different correction. Suppose the individual microstate is not observed, e.g. as in the case of black hole, then rather than trying to determine the splitting for each individual states, it is more meaningful to compute some kind of average correction over all microstates of the level $n=0$. As the system is in isolation, we have a microcanonical ensemble where probability of occurance is equally likely for all the microstates, i.e. $p_{i}=1/\Omega_{0}$. As a result, what is observed physically is the ensemble average
$\langle h^{\prime}\rangle:=\frac{1}{\Omega_{0}}\sum_{|i\rangle\in\mathbb{V}_{0}}\langle i|h^{\prime}|i\rangle.$ (A.11)
In the main text of the paper, we have claimed that $\langle h^{\prime}\rangle=0$ for the displaced fuzzy spheres system. To see this, let us use
@@
$h^{\prime}=P^{pq}\xi^{p\dagger}_{k}\xi^{q}_{k}-Q^{pq}\chi^{p\dagger}_{k}\chi^{q}_{k}+T^{pq}\xi^{p\dagger}_{k}\chi^{q\dagger}_{k}+{\rm h.c.}$ (A.13)
@@
$P^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal U}^{q}_{m\beta}\tilde{h}^{\prime}_{\alpha\beta},\quad Q^{pq}:={\cal V}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\tilde{h}^{\prime}_{\alpha\beta},\quad$ $T^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\tilde{h}^{\prime}_{\alpha\beta}.$ (A.14)
It is easy to see that the $T$ terms do not contribute to the matrix element. Also, we have the matrix elements
$\langle i|\xi^{p\dagger}_{k}\xi^{q}_{k}|i\rangle=n^{\xi}_{i}\delta^{pq},\quad\langle i|\chi^{p\dagger}_{k}\chi^{q}_{k}|i\rangle=n^{\chi}_{i}\delta^{pq},$ (A.15)
where
$n^{\xi}_{i}=N\binom{N^{2}-1}{r}\binom{N^{2}}{s},\quad n^{\chi}_{i}=N\binom{N^{2}-1}{s}\binom{N^{2}}{r}$ (A.16)
for a state $|i\rangle$ of the form (A.7). Therefore
$\langle i|h^{\prime}|i\rangle=n^{\xi}_{i}\mbox{tr}P-n^{\chi}_{i}\mbox{tr}Q.$ (A.17)
Using $\sum_{r}\binom{N^{2}-1}{r}\binom{N^{2}}{N^{2}-r}=\Omega_{0}/2$, the sum over all states of form (A.7) with a fixed level $n=0$ is given by
@@
[…]
where $K=\sigma^{a}X_{0}^{a}$ is the kernel for the unperturbed Hamiltonian $h_{0}$. We emphasis that the result (A.19) holds true in general and is independent of the form of $\tilde{h}^{\prime}$. In this paper, we consider a perturbation with the kernel
$\tilde{h}^{\prime}=\sigma^{a}x^{a}$ (A.20)
due to a separation of the fuzzy spheres. As a result, $\langle h^{\prime}\rangle=0$ since $\mbox{tr}(X^{0a})=0$. We note that this result is also expected since the perturbation (A.20) takes the form of a spin in an external magnetic field $\vec{x}$. As the fuzzy sphere on which the microstates are defined is isotropic, the magnetic energy averaged over the set of microstates is zero.